# Knowledge Transfer-based Evolutionary Deep Neural Network for Intelligent Fault Diagnosis

### v4

@@ Abstract
Genetic Algorithm , Multi-objective Optimization , Non-dominated sorting genetic algorithm (NSGA) II , Intelligent Fault Diagnosis , Knowledge transfer , Automatic Architecture Search
journal: Computers & Electrical Engineering
[…]
[inst1]organization=Dept. of Electrical Engineering, Indian Institute of Technology,city=Kanpur, postcode=208016, state=U.P., country=India
1 Introduction
@@
With the increase in complexity in modern industrial machinery, continuous monitoring has become a major concern to ensure the reliability, smooth operation, and productivity of systems. Therefore, continuous fault diagnosis, also called condition-based monitoring (CBM), is an essential requirement of today’s industries and has gained much attention from researchers worldwide [1, 2, 3, 4, 5, 6]. Progress in the area of fault diagnosis of industrial machines can be divided roughly into three stages: (i) traditional methods, (ii) fault diagnosis based on signal processing and analysis, and (iii) intelligent fault diagnosis.
@@
Traditional fault diagnosis methods mainly rely on experience with machine running conditions. The signal processing and analysis-based fault diagnosis requires complex mathematical analysis to identify specific changes in the machine signals like current, vibration, temperature, acoustic transmissions, etc. [7, 8, 9]. Therefore, these methods are not suitable for continuous monitoring in modern industrial scenarios. In recent decades, intelligent fault diagnosis has been the most investigated method due to the learning capability of machine learning models [10, 11, 12, 13, 14, 15]. Support vector machine (SVM) [11] has been reported with outstanding performance for intelligent pattern recognition. However, it fails to perform well for big data classification with high sparsity. For such problems, applications of deep neural networks (DNN) have been reported to be very effective [12, 13]. DNN is highly capable of hierarchical feature transformation of raw data to linearly separable features. Recently, several feature extraction methods using genetic algorithms have been reported, [16, 17, 18] for the training of DNN. However, for continuous monitoring under changeable machine working conditions, training a new DNN for the new dataset is an uneconomical and time-consuming task.
@@
The concept of transfer learning accelerated the DNN training process for the new data domain by initializing the weight parameters using knowledge of the pre-trained model on the source data [19, 20, 21]. Further, to solve the problem of domain shift, domain adaptation based on transfer component analysis was introduced in 2011 by S.J. Pan [22]. Later, the concept of domain adaptation using minimization of Maximum Mean Discrepancy (MMD) gained much attention by various researchers [23, 24, 25, 26, 27, 28, 29]. Long et. al. [23] used the principle of structural risk minimization and regularization theory to obtain a mechanism of adaptation regularization for cross-domain learning. L. Wen et. al. [26] suggested a fine-tuning method by minimizing the classification loss and the MMD term calculated for labeled source data and the labeled target data. The performance of all these methods is greatly affected by the selection of model architecture. Therefore, fault diagnosis under variable operating conditions of machines demands a fast network architecture search method capable of finding the best suitable model for the given working condition using a small amount of leveled target data.
@@
The motivation behind this work is to develop an automatic architecture optimization algorithm for DNN to classify accurately in the new data domain. Various methods for the neural architecture search have been reported in literature [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 14, 41, 42]. [30, 31] suggested the use of random search and greedy sequential algorithm for the optimization of hyperparameters of neural networks and deep belief networks. Random search or sequential search algorithms explore all possible architectures in a given search space. Therefore, exploring for the best suitable architecture takes too long. Baker et. al. [32] introduced a guided search algorithm using a Q-learning agent for the automatic selection of CNN architecture. In recent decades, genetic algorithms (GA) have gained much attention for hyperparameter optimization due to their multi-constraint optimization capabilities [39, 37, 38]. However, the application of GA to deep learning algorithms is a complex problem due to a large number of constraint variables and a time-consuming fitness evaluation (training and testing) process. Further, Luo et. al. [14] proposed Wasserstein GAN meta-learning (WGANML) that uses the concept of the game between the generative and discriminative models to generate missing samples, followed by meta-learning for parameter adjustment for better results. Sun et. al. [38] designed a variable-length gene encoding technique to encode the CNN architecture and adopted the fitness evaluation strategy of training a CNN model for a few iterations to find the best of the generation. Each individual (CNN model) in the population is initialized using Gaussian distribution and trained from scratch. This method requires the training of CNN models from scratch at each generation. Therefore, the evolution of architecture for the best performance becomes a time-consuming process.
@@
The organization of the rest of the paper is as follows. Section 2 defines the problem objective of automatic architecture search. Section 3 briefly introduces the theoretical background of evolutionary algorithms, multi-objective optimization, DNN, and knowledge transfer from a DNN to another through the function-preserving principle. Section 4 describes the proposed EvoN2N. Section 5 discusses the effectiveness of the proposed framework and its comparison with state-of-the-art methods on (i) Case Western Reserve University (CWRU) dataset [45], (ii) the Paderborn University (PBU) dataset [46] and (iii) the gearbox fault detection dataset [47] under different operating conditions. Finally, Section 6 concludes the work.
2 Problem Statement
Let $W^{s}$ be the teacher (source) model trained on a dataset $\mathcal{D}^{s}=\left(X^{s},Y^{s}\right)$ from the source domain and $\mathcal{D}^{tr}=\left(X^{tr},Y^{tr}\right)$, $\mathcal{D}^{val}=\left(X^{val},Y^{val}\right)$, & $\mathcal{D}^{te}=\left(X^{te},Y^{te}\right)$ are the training dataset, the validation dataset, & the test dataset respectively from the target dataset $\mathcal{D}^{t}=\left(X^{t},Y^{t}\right)$. The objective of fault diagnosis via automatic DNN architecture search can mathematically be expressed as
@@ 2 Problem Statement
3 Theoretical Background
3.1 Multi-objective Genetic Algorithm
Multi-Objective Genetic Algorithm (MOGA) is a heuristic search-based optimization technique under multi-constraint to find multiple optimal solutions called Pareto-optimal solution. A set of multiple solutions is obtained, out of which, any one solution cannot be said to be better without some additional information. For such problems, several MOGAs have been reported in various publications [48, 49, 50, 51, 52, 53, 54]. The non-dominated sorting genetic algorithm (NSGA-II) [52] has gained much attention due to its fast sorting methodology. Our problem consists of unconstrained optimization of the depth & width of a DNN architecture for maximum diagnostic performance while keeping the number of total trainable parameters minimal. Therefore, we have considered architecture optimization as a multi-objective optimization problem instead of a single-objective optimization problem. We have selected the NSGA-II framework for the evolution of DNN architecture, considering the objectives (i) to maximize the validation accuracy and (ii) to minimize the total number of weight matrices.
3.2 Deep Neural Network (DNN)
A DNN is a multi-layered neural network formed by stacked auto-encoder (SAE) [55] with softmax classifier as output layer. DNN has the capability of highly non-linear function approximation. Each layer of SAE is trained by a greedy layer unsupervised learning mechanism followed by stacking together to form SAE. Now, SAE with softmax classifier at the output layer is fine-tuned by gradient descent using a labeled dataset. Fig. 1 shows a general structure of DNN with softmax classifier as the output layer.
Figure 1: General structure of multi-layered DNN with softmax classifier: $h_{i}$ represents the number of nodes in the $i^{th}$ hidden layer and $W_{ji}$ represents the weight matrix connecting $i{th}$ to $j^{th}$ layer.
3.3 Domain Adaptation
If the target datasets have different distributions from the source data, the DNN trained on the source data fails to classify correctly on the target data. [25]. The diagnostic model trained on the source data needs to be fine-tuned using the target data to shift the classifier shown in Fig. 2.
4 Proposed Methodology
@@
7:$Q\longleftarrow\textbf{CrossOverMutation}(P)$   //Apply crossover and mutation (Section 4.4).
@@
15:$Q\longleftarrow\textbf{CrossOverMutation}(P)$   //Apply crossover and mutation (Section 4.4).
@@
4.1 Gene Encoding and Population Initialization
@@ 4.2 Evaluate Fitness
4.3 Parent Selection using Rank and Crowded Distance
@@ 4.4 Crossover and Mutation
5 Experimental Results and Discussion
@@
5.1 Setup Description
5.1.1 CWRU Bearing Data [45]
@@ 5.1.2 PBU Dataset [46]
5.1.3 Gearbox Fault Detection (GFD) Dataset [47]
@@
5.2 Data Processing and Evaluation Scheme
@@
    Source Data: The time-series signals recorded at 12 Hz DE with a motor load of 00 hp and speed of $1796$ RPM are segmented to create $1200\times 100$ samples per class. After merging the segmented dataset from each class (N, IR, OR, and B), the it contains a total of $4800\times 100$ samples.
@@
    Target-1: The time-series signals recorded at 12k Hz FE with motor loads of $1$, $2$, & $3$ hp, fault diameter (F.D.) varying from $7$ to $21$ mil, and speed of $1772$ RPM are used to prepare the dataset for target-1 (T1) under nine different cases with F.D.’s and motor loads as shown in table 1.
@@
    Target-2 & 3: Two different sets of target datasets (target-2 (T2) & target-3 (T3)) are created by considering different levels of damage (extent of damage) under four L.S.’s of LS1, LS2, LS3, & LS4 as given in table 2.
@@
    Target-4: Target-4 (T4) is created under three different load conditions using the signals from the files corresponding to 30, 60, &90 % of load as summarised in Table 3. For each load, a small number of sample points $30000$ are used to create a segmented data set of $300\times 100$ samples per class.
@@
Table 1: CA for CWRU bearing fault data (T1) recorded at Fan End (FE) with three fault diameters (F.D.s)
@@
Table 2: CA for PBU Dataset (T2 & T3) with different levels of fault and different load settings (L.S.)
@@
Table 3: CA for Gearbox Fault Detection Dataset (T4) with 30, 50, 70, 90% of load
@@
5.3 Evaluation metrics
Classification performance of a diagnostic model is measured in terms of classification precision ($CA$) as widely accepted in the literature [23, 25, 26].
@@ 5.3 Evaluation metrics
5.4 Implementation and Training Parameters
The proposed EvoN2N as described in Algorithm 1 requires an initial DNN model ($\Psi^{s}$) termed as the teacher model or the source model. Therefore, first, a source model with architecture (we choose a DNN model with $3$ hidden layers $80-40-20$) is trained on the source dataset (as mentioned in Section 5.2 for three cases) using back-propagation gradient descent.
@@ 5.4 Implementation and Training Parameters
The performance of the proposed framework is compared with the state-of-the-art method most popularly used for intelligent fault diagnosis. The selected algorithms are support vector machines (SVM) [11], deep neural network (DNN) [12], deep transfer learning (DTL) based on a sparse autoencoder [26], Deep neural network for domain Adaptation in Fault Diagnosis (DAFD) [25], Net2Net without domain adaptation (N2N_WDA) [28], Net2Net with domain adaptation (N2N_DA) [28], evolutionary deep CNN (EvoDCNN) [39], and Wasserstein GAN meta-learning (WGANML) [14]. The architecture for DNN, DTL, and DAFD is kept the same as the source network. The population size and the maximum number of generations for EvoDCNN are kept the same as the proposed method. All these models are trained on the same data as described in section 5.2 using the methods suggested in the cited references. The classification accuracies are tabulated in tables 1 and 2. For better visualization of the classification performance, the confusion matrix for one of the cases taken from T1 (7mil, 1hp) has been shown in Fig. 7. For comparative analysis, Fig. 8 shows the transfer improvement ($TIs$) in terms of average $CA$ ($\overline{CA}$) for the three cases: (i) CWRU dataset, (ii) PBU dataset, and (iii) GFD dataset.
@@ 5.5 Ablation Study:
5.5.1 Effect of N2N-based quick learning
@@
5.5.2 Effect of population size
Table 4. shows the optimal architecture obtained at various generations for the population size of 50 and 100. It can be seen that the proposed framework does not converse fully in 20 generations with a population size of 50, however, it converses faster with the larger population size.
Table 4: Evolution of best architecture at every $5^{th}$ generation (CWRU FE, 7 mil FD, 1 hp load)
@@ 5.5.2 Effect of population size
5.5.3 Effect of number of generation
From Table 4 and the accuracy curve shown in Fig. 6, it can be observed that the algorithm takes about 20 generations to reach to its optimal for each of the dataset cases.
5.6 Hypothesis Testing
@@
Table 5: Rank table for results shown in Tables 1, 2, and 3 for all datasets T1-T4.
@@
For the hypothesis test, a combined rank table for results shown in Tables 1, 2, and 3 is created, which constitute a total of 21 sub-cases of datasets and 9 classifiers (methods). The average ranks of all the algorithms for the 21 datasets are calculated and shown in Table 5.
For the rank table shown in Table 5, Friedman’s ANOVA table is obtained using MATLAB’s function ‘friedman’, and the ${p}$-value is compared on the significance level of 0.05 and 0.01. Friedman’s ANOVA table is shown in Table 6.
Table 6: Friedman’s ANOVA Table: $S_{S}$ = Sum of squares and $d_{F}$ = degree of freedom.
@@
5.7 Discussion
Following points can be observed from the diagnostic results shown in tables 1, 2, & 3 and Fig.’s 7, 8, & 6.

### v5

@@ Abstract
Intelligent Fault Diagnosis, Multi-objective Optimization, Knowledge Transfer, Automatic Architecture Search,
I Introduction
@@
With the increase in complexity in modern industrial machinery, continuous monitoring has become a major concern to ensure the reliability, smooth operation, and productivity of systems. Therefore, continuous fault diagnosis, also called condition-based monitoring (CBM), is an essential requirement of today’s industries and has gained much attention from researchers worldwide [1, 2, 3]. Progress in the area of fault diagnosis of industrial machines can be divided roughly into three stages: (i) traditional methods, (ii) fault diagnosis based on signal processing and analysis, and (iii) intelligent fault diagnosis.
@@
Traditional fault diagnosis methods mainly rely on experience with machine running conditions. The signal processing and analysis-based fault diagnosis requires complex mathematical analysis to identify specific changes in the machine signals like current, vibration, temperature, acoustic transmissions, etc. [4, 5]. Therefore, these methods are not suitable for continuous monitoring in modern industrial scenarios. In recent decades, intelligent fault diagnosis has been the most investigated method due to the learning capability of machine learning models [6, 7, 8, 9, 10, 11]. Support vector machine (SVM) [7] has been reported with outstanding performance for intelligent pattern recognition. However, it fails to perform well for big data classification with high sparsity. For such problems, applications of deep neural networks (DNN) have been reported to be very effective [8, 9]. DNN is highly capable of hierarchical feature transformation of raw data to linearly separable features. Recently, several feature extraction methods using genetic algorithms have been reported, [12, 13, 14] for the training of DNN. However, for continuous monitoring under changeable machine working conditions, training a new DNN for the new dataset is an uneconomical and time-consuming task.
@@
The concept of transfer learning accelerated the DNN training process for the new data domain by initializing the weight parameters using knowledge of the pre-trained model on the source data [15, 16]. Further, to solve the problem of domain shift, domain adaptation based on transfer component analysis was introduced in 2011 by S.J. Pan [17]. Later, the concept of domain adaptation using minimization of Maximum Mean Discrepancy (MMD) gained much attention by various researchers [18, 19, 20, 21, 22]. Long et. al. [18] used the principle of structural risk minimization and regularization theory to obtain a mechanism of adaptation regularization for cross-domain learning. L. Wen et. al. [19] suggested a fine-tuning method by minimizing the classification loss and the MMD term calculated for labeled source data and the labeled target data. The performance of all these methods is greatly affected by the selection of model architecture. Therefore, fault diagnosis under variable operating conditions of machines demands a fast network architecture search method capable of finding the best suitable model for the given working condition using a small amount of leveled target data.
@@
The motivation behind this work is to develop an automatic architecture optimization algorithm for DNN to classify accurately in the new data domain. Various methods for the neural architecture search have been reported in literature [23, 24, 25, 26, 27, 28, 29, 10, 30]. [23, 24] suggested the use of random search and greedy sequential algorithm for the optimization of hyperparameters of neural networks and deep belief networks. Random search or sequential search algorithms explore all possible architectures in a given search space. Therefore, exploring for the best suitable architecture takes too long. Baker et. al. [25] introduced a guided search algorithm using a Q-learning agent for the automatic selection of CNN architecture. In recent decades, genetic algorithms (GA) have gained much attention for hyperparameter optimization due to their multi-constraint optimization capabilities [29, 27, 28]. However, the application of GA to deep learning algorithms is a complex problem due to a large number of constraint variables and a time-consuming fitness evaluation (training and testing) process. Further, Luo et. al. [10] proposed Wasserstein GAN meta-learning (WGANML) that uses the concept of the game between the generative and discriminative models to generate missing samples, followed by meta-learning for parameter adjustment for better results. Sun et. al. [28] designed a variable-length gene encoding technique to encode the CNN architecture and adopted the fitness evaluation strategy of training a CNN model for a few iterations to find the best of the generation. Each individual (CNN model) in the population is initialized using Gaussian distribution and trained from scratch. This method requires the training of CNN models from scratch at each generation. Therefore, the evolution of architecture for the best performance becomes a time-consuming process.
@@
The organization of the rest of the paper is as follows. Section III-A defines the problem objective of automatic architecture search. Section II briefly introduces the theoretical background of evolutionary algorithms, multi-objective optimization, DNN, and knowledge transfer from a DNN to another through the function-preserving principle. Section III-B describes the proposed EvoN2N. Section IV discusses the effectiveness of the proposed framework and its comparison with state-of-the-art methods on (i) Case Western Reserve University (CWRU) dataset [33], (ii) the Paderborn University (PBU) dataset [34] and (iii) the gearbox fault detection dataset [35] under different operating conditions. Finally, Section V concludes the work.
II Theoretical Background
II-A Multi-objective Genetic Algorithm
Multi-Objective Genetic Algorithm (MOGA) is a heuristic search-based optimization technique under multi-constraint to find multiple optimal solutions called Pareto-optimal solution. A set of multiple solutions is obtained, out of which, any one solution cannot be said to be better without some additional information. For such problems, several MOGAs have been reported in various publications [36, 37, 38, 39]. The non-dominated sorting genetic algorithm (NSGA-II) [37] has gained much attention due to its fast sorting methodology. Our problem consists of unconstrained optimization of the depth & width of a DNN architecture for maximum diagnostic performance while keeping the number of total trainable parameters minimal. Therefore, we have considered architecture optimization as a multi-objective optimization problem instead of a single-objective optimization problem. We have selected the NSGA-II framework for the evolution of DNN architecture, considering the objectives (i) to maximize the validation accuracy and (ii) to minimize the total number of weight matrices.
II-B Domain Adaptation
If the target datasets have different distributions from the source data, the DNN trained on the source data fails to classify correctly on the target data. The diagnostic model trained on the source data needs to be fine-tuned using the target data to shift the classifier shown in Fig. 1.
III Proposed Methodology
III-A Problem Statement
A DNN is a multi-layered neural network formed by stacked auto-encoder (SAE) [41] with softmax classifier as output layer. DNN has the capability of highly non-linear function approximation. Each layer of SAE is trained by a greedy layer unsupervised learning mechanism followed by stacking together to form SAE. Now, SAE with softmax classifier at the output layer is fine-tuned by gradient descent using a labeled dataset. Let $W^{s}$ be the teacher (source) model trained on a dataset $\mathcal{D}^{s}=\left(X^{s},Y^{s}\right)$ from the source domain and $\mathcal{D}^{tr}=\left(X^{tr},Y^{tr}\right)$, $\mathcal{D}^{val}=\left(X^{val},Y^{val}\right)$, & $\mathcal{D}^{te}=\left(X^{te},Y^{te}\right)$ are the training dataset, the validation dataset, & the test dataset respectively from the target dataset $\mathcal{D}^{t}=\left(X^{t},Y^{t}\right)$. The objective of fault diagnosis via automatic DNN architecture search can mathematically be expressed as
@@ III-A Problem Statement
III-B Proposed Framework
@@
7:$Q\longleftarrow\textbf{CrossOverMutation}(P)$   //Apply crossover and mutation (Section III-B4).
@@
15:$Q\longleftarrow\textbf{CrossOverMutation}(P)$   //Apply crossover and mutation (Section III-B4).
@@
III-B1 Gene Encoding and Population Initialization
@@ III-B2 Fitness Evaluation
III-B3 Parent Selection using Rank and Crowded Distance
@@ III-B4 Crossover and Mutation
IV Experimental Results and Discussion
@@
IV-A Setup Description
IV-A1 CWRU Bearing Data [33]
@@ IV-A2 PBU Dataset [34]
IV-A3 Gearbox Fault Detection (GFD) Dataset [35]
@@
IV-B Data Processing and Evaluation Scheme
@@
    Source Data: The time-series signals recorded at 12 Hz DE with a motor load of 00 hp and speed of $1796$ RPM are segmented to create $1200\times 100$ samples per class. After merging the segmented dataset from each class (N, IR, OR, and B), it contains a total of $4800\times 100$ samples.
@@
    Target-1: The time-series signals recorded at 12k Hz FE with motor loads of $1$, $2$, & $3$ hp, fault diameter (F.D.) varying from $7$ to $21$ mil, and speed of $1772$ RPM are used to prepare the dataset for target-1 (T1) under nine different cases with F.D.’s and motor loads as shown in table I.
@@
    Target-2 & 3: Two different sets of target datasets (target-2 (T2) & target-3 (T3)) are created by considering different levels of damage (extent of damage) under four L.S.’s of LS1, LS2, LS3, & LS4 as given in table II.
@@
    Target-4: Target-4 (T4) is created under three different load conditions using the signals from the files corresponding to 30, 60, &90 % of load as summarised in Table III. For each load, a small number of sample points $30000$ are used to create a segmented data set of $300\times 100$ samples per class.
@@
Table I: CA for CWRU bearing fault data (T1) recorded at Fan End (FE) with three fault diameters (F.D.s)
@@
Table II: CA for PBU Dataset (T2 & T3) with different levels of fault and different load settings (L.S.)
@@
Table III: CA for Gearbox Fault Detection Dataset (T4) with 30, 50, 70, 90% of load
@@
IV-C Evaluation metrics
Classification performance of a diagnostic model is measured in terms of classification precision ($CA$) as widely accepted in the literature [18, 19].
@@ IV-C Evaluation metrics
IV-D Implementation and Training Parameters
The proposed EvoN2N as described in Algorithm 1 requires an initial DNN model ($\Psi^{s}$) termed as the teacher model or the source model. Therefore, first, a source model with architecture (we choose a DNN model with $3$ hidden layers $80-40-20$) is trained on the source dataset (as mentioned in Section IV-B for three cases) using back-propagation gradient descent.
@@ IV-D Implementation and Training Parameters
The performance of the proposed framework is compared with the state-of-the-art method most popularly used for intelligent fault diagnosis. The selected algorithms are support vector machines (SVM): the baseline method [7], deep neural network (DNN) [8], Domain Adversarial Graph Convolutional Network (DAGCN) [22], Net2Net without domain adaptation (N2N_WDA) [21], Net2Net with domain adaptation (N2N_DA) [21], evolutionary deep CNN (EvoDCNN) [29], DenseNAS [26], and Wasserstein GAN meta-learning (WGANML) [10]. The architecture for DNN is kept the same as the source network. The model architectures and hyperparameters for other models are chosen based on standard values provided in their references. The population size and the maximum number of generations for EvoDCNN are kept the same as the proposed method. All these models are trained on the same data as described in section IV-B using the methods suggested in the cited references. The classification accuracies are tabulated in tables I and II. For better visualization of the classification performance, the confusion matrix for one of the cases taken from T1 (7mil, 1hp) has been shown in Fig. 6. For comparative analysis, Fig. 7 shows the transfer improvement ($TIs$) in terms of average $CA$ ($\overline{CA}$) for the three cases: (i) CWRU dataset, (ii) PBU dataset, and (iii) GFD dataset.
@@ IV-E Ablation Study:
IV-E1 Effect of N2N-based quick learning
@@
IV-E2 Effect of population size
Table IV. shows the optimal architecture obtained at various generations for the population size of 50 and 100. It can be seen that the proposed framework does not converse fully in 20 generations with a population size of 50, however, it converses faster with the larger population size.
Table IV: Evolution of best architecture at every $5^{th}$ generation (CWRU FE, 7 mil FD, 1 hp load)
@@ IV-E2 Effect of population size
IV-E3 Effect of number of generation
From Table IV and the accuracy curve shown in Fig. 5, it can be observed that the algorithm takes about 20 generations to reach to its optimal for each of the dataset cases.
IV-F Hypothesis Testing
@@
Table V: Rank table for results shown in Tables I, II, and III for all datasets T1-T4.
@@
For the hypothesis test, a combined rank table for results shown in Tables I, II, and III is created, which constitute a total of 21 sub-cases of datasets and 9 classifiers (methods). The average ranks of all the algorithms for the 21 datasets are calculated and shown in Table V.
For the rank table shown in Table V, Friedman’s ANOVA table is obtained using MATLAB’s function ‘friedman’, and the ${p}$-value is compared on the significance level of 0.05 and 0.01. Friedman’s ANOVA table is shown in Table VI.
Table VI: Friedman’s ANOVA Table: $S_{S}$ = Sum of squares and $d_{F}$ = degree of freedom.
@@
IV-G Discussion
Following points can be observed from the diagnostic results shown in tables I, II, & III and Fig.’s 6, 7, & 5.